# Failure of famous functional inequalities on Finsler manifolds: the influence of $S$-curvature

### v1

@@ 6.2. Counterexamples under $\mathbf{Ric}_{N}\geq-K^{2}$
Our main results are valid under the separate curvature assumptions
$\mathbf{Ric}\geq-(n-1)k^{2},\quad\mathbf{S}\geq(n-1){h}.$ (6.13)
Recall that there is another kind of Ricci curvature on such a space, which is a combination of $\mathbf{Ric}$ and $\mathbf{S}$, see Ohta [21]. More precisely, for a $\operatorname{\mathsf{FMMM}}$ $(M,F,\mathfrak{m})$, the weighted Ricci curvature $\mathbf{Ric}_{N}$ is defined as follows: given $N\in[n,\infty]$, for any unit vector $y\in SM$,
$\mathbf{Ric}_{N}(y)=\left\{\begin{array}[]{lll}\mathbf{Ric}(y)+\left.\frac{\operatorname{d}}{{\operatorname{d}}t}\right|_{t=0}\mathbf{S}(\dot{\gamma}_{y}(t))-\frac{\mathbf{S}^{2}(y)}{N-n},&&\text{ for }N\in(n,\infty),\\ \\ \underset{L\downarrow n}{\lim}\mathbf{Ric}_{L}(y),&&\text{ for }N=n,\\ \\ \mathbf{Ric}(y)+\left.\frac{\operatorname{d}}{{\operatorname{d}}t}\right|_{t=0}\mathbf{S}(\dot{\gamma}_{y}(t)),&&\text{ for }N=\infty.\end{array}\right.$ (6.19)
Weighted Ricci curvature is the realization of $\operatorname{\mathsf{CD}}(K,N)$ condition in the Finsler framework, which is a Finslerian version of the Bakry–Émery Ricci tensor. In view of this notion, it is a natural question whether or not the separated condition (6.13) can be replaced by $\mathbf{Ric}_{N}\geq-K^{2}$ for some $K\geq 0$. In fact, the following examples show that the statements of Theorems 1.1 &\&& 1.2 fail under $\mathbf{Ric}_{N}\geq-K^{2}$.
@@
(see Li [19]) There are infinitely many reversible Minkowski norms on Euclidean spaces. In the sequel, we provide two such classes:
@@
    $\phi(y)=\sqrt[2m]{\sum^{n}_{i=1}(y^{i})^{2m}},\ y=(y_{1},...,y_{n})\in\mathbb{R}^{n},$
@@
    Let $|y|$ and $|\tilde{y}|$ be the Euclidean norms on $\mathbb{R}^{n}$ and $\mathbb{R}^{m}$, respectively. Then
    $\phi(y,\tilde{y})=\sqrt{\sqrt{|y|^{4}+|\tilde{y}|^{4}}+|y|^{2}+|\tilde{y}|^{2}},\ (y,\tilde{y})\in\mathbb{R}^{n}\times\mathbb{R}^{m},$
@@
In general, given a reversible Minkowski norm $\phi$ on the Euclidean space $\mathbb{R}^{n}$, the triple $(\mathbb{R}^{n},\phi,\mathscr{L}^{n})$ is a complete noncompact $\operatorname{\mathsf{FMMM}}$ with $\mathbf{Ric}_{N}\equiv 0$ for all $n\in[N,\infty]$. However, according to Huang, Kristaly and Zhao [10], the $L^{p}$-Hardy inequality, Heisenberg–Pauli–Weyl uncertainty principle and Caffarelli–Kohn–Nirenberg interpolation inequality are valid as well as sharp with the same optimal constants as in the Euclidean case; we also notice that $\mathbf{S}\equiv 0$ for such spaces.
@@
We consider the triplet $(\Omega,\tilde{F},\operatorname{\mathfrak{m}}_{BH})$, where $\operatorname{\mathfrak{m}}_{BH}$ is the Busemann–Hausdorff measure of $F$ (not $\tilde{F}$). We first observe that $\tilde{F}$ is reversible, i.e., $\lambda_{\tilde{F}}(\Omega)=1$. Then, according to Shen [26], $(\Omega,\tilde{F},\operatorname{\mathfrak{m}}_{BH})$ is a complete, $n$-dimensional projectively flat $\operatorname{\mathsf{FMMM}}$ of constant flag curvature $\widetilde{\mathbf{K}}=-1$. In addition, due to Ohta [22], one has that
@@
These properties clearly show that in Theorem 1.2 we cannot replace the combined curvature conditions with the lower boundedness of the weighted Ricci curvature.
@@
First, we observe that $\widetilde{\mathbf{Ric}}=-(n-1)$, as $\widetilde{\mathbf{K}}=-1$. Moreover, one has that
$-(n+1)\leq\tilde{\mathbf{S}}\leq n+1.$ (6.22)
Indeed, according to Okada [23], one has that every Funk metric satisfies
@@
Let $P=P(x,y)$ denote the projective factor of the Funk metric $F$. Thus, (6.23) yields
@@
Consider the Busemann–Hausdorff measure $\operatorname{\mathfrak{m}}_{BH}$ of $F$, whose density function $\sigma(x)$ is a constant, see (6.11). Since $\tilde{G}^{m}=\tilde{P}y^{m}$, relation (2.11) yields
@@
Finally, according to Troyanov [32, p. 103], we have
@@
Thus, for any $y\in\widetilde{S\Omega}:=\{v\in T\Omega\,:\,\tilde{F}(v)=1\}$, by (2.6), (6.25) and (6.26), we obtain
$\left.\frac{\operatorname{d}}{{\operatorname{d}}t}\right|_{t=0}\tilde{\mathbf{S}}(\dot{\gamma}_{y}(t))$ $=\frac{\partial\tilde{\mathbf{S}}}{\partial x^{k}}y^{k}-2\frac{\partial\tilde{\mathbf{S}}}{\partial y^{m}}\tilde{G}^{m}(y)=(n+1)\left(\frac{\partial\tilde{P}}{\partial x^{k}}y^{k}-2\frac{\partial\tilde{{P}}}{\partial y^{m}}\tilde{G}^{m}(y)\right)$

### v2

@@ 6.2. Counterexamples under $\mathbf{Ric}_{N}\geq-K^{2}$
This subsection demonstrates the necessity of our separate curvature assumptions
$\mathbf{Ric}\geq-(n-1)k^{2},\quad\mathbf{S}\geq(n-1){h}$ (6.13)
by showing that the main results fail if those are replaced by a lower bound on the weighted Ricci curvature $\mathbf{Ric}_{N}$.
Recall that for a Finsler metric measure manifold $(M,F,\mathfrak{m})$, the weighted Ricci curvature $\mathbf{Ric}_{N}$ combines both $\mathbf{Ric}$ and $\mathbf{S}$, which is a Finslerian version of the Bakry–Émery Ricci tensor. The lower bound $\mathbf{Ric}_{N}\geq K$ is equivalent to the $\mathrm{CD}(K,N)$ condition (see Ohta [21, 22, 23]).
More precisely, the weighted Ricci curvature $\mathbf{Ric}_{N}$ is defined as follows: given $N\in[n,\infty]$, for any unit vector $y\in SM$,
$\mathbf{Ric}_{N}(y)=\left\{\begin{array}[]{lll}\mathbf{Ric}(y)+\left.\frac{\operatorname{d}}{{\operatorname{d}}t}\right|_{t=0}\mathbf{S}(\gamma^{\prime}_{y}(t))-\frac{\mathbf{S}^{2}(y)}{N-n},&&\text{ for }N\in(n,\infty),\\ \\ \underset{L\downarrow n}{\lim}\mathbf{Ric}_{L}(y),&&\text{ for }N=n,\\ \\ \mathbf{Ric}(y)+\left.\frac{\operatorname{d}}{{\operatorname{d}}t}\right|_{t=0}\mathbf{S}(\gamma^{\prime}_{y}(t)),&&\text{ for }N=\infty.\end{array}\right$. (6.19)
Although it seems natural to replace the separated condition (6.13) by $\mathbf{Ric}_{N}\geq-K^{2}$ for some $K\geq 0$, the following examples demonstrate that Theorems 1.1 and 1.2 will fail.
@@
(Li [19]) There are infinitely many reversible Minkowski norms on Euclidean spaces. In the sequel, we provide two such classes:
@@
    $\phi(y)=\sqrt[2m]{\sum^{n}_{i=1}(y^{i})^{2m}},\quad\forall\,y=(y_{1},...,y_{n})\in\mathbb{R}^{n},$
@@
[…]
    $\phi(y,\tilde{y})=\sqrt{\sqrt{|y|^{4}+|\tilde{y}|^{4}}+|y|^{2}+|\tilde{y}|^{2}},\quad\forall\,(y,\tilde{y})\in\mathbb{R}^{n}\times\mathbb{R}^{m},$
@@
In general, given a reversible Minkowski norm $\phi$ on the Euclidean space $\mathbb{R}^{n}$, the triple $(\mathbb{R}^{n},\phi,\mathscr{L}^{n})$ is a complete noncompact $\operatorname{\mathsf{FMMM}}$ with $\mathbf{Ric}_{N}\equiv 0$ for all $n\in[N,\infty]$. However, according to [10], the $L^{p}$-Hardy inequality, Heisenberg–Pauli–Weyl uncertainty principle and Caffarelli–Kohn–Nirenberg interpolation inequality are valid as well as sharp with the same optimal constants as in the Euclidean case; we also notice that $\mathbf{S}\equiv 0$ for such spaces.
@@
Consider the triplet $(\Omega,\tilde{F},\operatorname{\mathfrak{m}}_{BH})$, where $\operatorname{\mathfrak{m}}_{BH}$ is the Busemann–Hausdorff measure of $F$ (not $\tilde{F}$), which is a constant multiple of Lebesgue measure (see (6.11)). Note that $\tilde{F}$ is reversible, i.e., $\lambda_{\tilde{F}}(\Omega)=1$. Then, by [27], $(\Omega,\tilde{F},\operatorname{\mathfrak{m}}_{BH})$ is an $n$-dimensional complete reversible $\operatorname{\mathsf{FMMM}}$ of constant flag curvature $\widetilde{\mathbf{K}}=-1$. Additionally, Ohta [23] shows that
@@
These properties demonstrate that in Theorem 1.2, the curvature conditions (6.13) cannot be replaced by a lower bound on the weighted Ricci curvature alone.
@@
First, we observe that $\widetilde{\mathbf{Ric}}=-(n-1)$, as $\widetilde{\mathbf{K}}=-1$. Moreover, we claim
$(n+1)\leq\tilde{\mathbf{S}}\leq n+1.$ (6.22)
Indeed, according to Okada [24], every Funk metric satisfies
@@
Let $P=P(x,y)$ denote the projective factor of the Funk metric $F$. Then (5.2) and (6.23) yield
@@
Consider the Busemann–Hausdorff measure $\operatorname{\mathfrak{m}}_{BH}$ of $F$, whose density function $\sigma(x)$ is a constant, see (6.11). Since $\tilde{G}^{m}=\tilde{P}y^{m}$, relations (2.11) and (5.3) yield
@@
Finally, according to Troyanov [33, p. 103], we have
@@
Then for any $y\in\widetilde{S\Omega}:=\{v\in T\Omega\,:\,\tilde{F}(v)=1\}$, by (2.6), (6.25), (5.3) and (6.26), we obtain
$\left.\frac{\operatorname{d}}{{\operatorname{d}}t}\right|_{t=0}\tilde{\mathbf{S}}({\gamma}^{\prime}_{y}(t))$ $=\frac{\partial\tilde{\mathbf{S}}}{\partial x^{k}}y^{k}-2\frac{\partial\tilde{\mathbf{S}}}{\partial y^{m}}\tilde{G}^{m}(y)=(n+1)\left(\frac{\partial\tilde{P}}{\partial x^{k}}y^{k}-2\frac{\partial\tilde{{P}}}{\partial y^{m}}\tilde{G}^{m}(y)\right)$